# What Is Missing in Multilingual Visual Reasoning and How to Fix It

### v1

@@ 7 Related Work
Gap between Proprietary GPT-4V and Open Models
Currently, there is a trend of shifting from openly releasing models to paid APIs Abdalla et al. (2023). Previous research on examining GPT-4V demonstrates its unprecedented multimodal capabilities, and there is still a gap between this proprietary model and other open source models Yang et al. (2023). However, it is still important for the community to have as strong open source multimodal models.
@@ 8 Conclusion
In conclusion, our study explores the evolving domain of multilingual visual reasoning. We observe a trend towards inclusivity in models, yet recognize persistent disparities in performance across languages and cultures. While proprietary systems like GPT-4V exhibit notable accuracy and equity across languages, open models still face challenges in bridging the gap, especially for low-resource languages. Our comparative analysis highlights the superior performance of GPT-4V but also underscores the need for advancements in open models. Through strategic interventions addressing multilinguality, multimodality, and reasoning, we showcase promising avenues for improvement. Leveraging our interventions, we demonstrate significant enhancements in open model performance, achieving state-of-the-art results under zero-shot settings for open models. Our findings emphasizes the potential for further advancements in language-equitable multimodal reasoning, with the aim of narrowing down the gap between human and machine performance, and the gap between proprietary and open models.
Acknowledgements:
The authors thank the members of CMU who gave useful feedback on an early version of the draft. This work was supported in part by grants from Google and DSTA Singapore.
@@ A.1.1 Zero-Shot Evaluation (no labeled data for task)
Large Language and Vision Assistant (LLaVA) is an open large multimodal model that is instruction tuned on machine-generated instruction-following data Liu et al. (2023c, a). LLaVA extends the capabilities of existing models by incorporating visual models and large language models. It connects a vision encoder CLIP and an LLM decoder. LLaVA is not explicitly trained to process multilingual data, but the LLM decoder (Vicuna is the default LLM) is known to have seen multilingual data in pretraining Chiang et al. (2023).

### v2

@@ 7 Related Work
Gap between Proprietary Models and Open Models
Currently, there is a trend of shifting from openly releasing models to paid APIs Abdalla et al. (2023). Previous research on examining GPT-4V and Gemini 1.5 Pro demonstrates its unprecedented multimodal capabilities, and there is still a gap between this proprietary model and other open source models Yang et al. (2023). However, it is still important for the community to have as strong open source multimodal models.
@@ 8 Conclusion
In conclusion, we explore the evolving domain of multilingual visual reasoning. We observe a trend towards inclusivity in models, yet recognize persistent disparities in performance across languages and cultures. While proprietary systems like GPT-4V exhibit notable and equitable accuracy across languages, open models still face challenges in bridging the gap, especially for low-resource languages. Our analysis highlights the superior performance of GPT-4V but also underscores the need for advancements in open models. Leveraging interventions addressing multilinguality, multimodality, and reasoning, we demonstrate significant enhancements in open model performance, achieving state-of-the-art results under zero-shot settings for open models. Our findings emphasizes the potential for further advancements in multilingual visual reasoning, with the aim of narrowing down the gap between human and machine performance, and the gap between proprietary and open models.
@@ A.1.1 Zero-Shot Evaluation (no labeled data for task)
Large Language and Vision Assistant (LLaVA) is a series of open large multimodal model that are instruction tuned on machine-generated instruction-following data Liu et al. (2023c, a, 2024). LLaVA extends the capabilities of existing models by incorporating visual models and large language models. It connects a vision encoder CLIP and an LLM decoder. LLaVA is not explicitly trained to process multilingual data, but the LLM decoder (Vicuna is the default LLM) is known to have seen multilingual data in pretraining Chiang et al. (2023).
Qwen-VL
Qwen-VL is an open large multilingual multimodal model trained on English and Chinese data. It is based on Qwen-7B Bai et al. (2023a), incorporating a language-aligned visual encoder and a positionaware adapter. It is trained to be able to process multi-image inputs.
Qwen2-VL
Following Qwen-VL, Qwen2-VL is also trained on English and Chinese data. It is based on Qwen2 Yang et al. (2024).
Cambrian
Cambrian-8B is a vision-centric multimodal LLM that focuses on bridging the gap between visual representation learning and language models Tong et al. (2024). It introduces the Spatial Vision Aggregator (SVA), which efficiently integrates high-resolution visual features with language models. Cambrian also offers a new benchmark called CV-Bench to evaluate 2D and 3D visual understanding. Through its open release of model weights, code, and datasets, Cambrian aims to foster advancements in multimodal AI systems and visual representation research.
Molmo
Trained from scratch, Molmo Deitke et al. (2024) is a family of models trained from scratch. It is especially trained on a special 2D-pointing dataset.
Llama3
Extending Llama 2 Touvron et al. (2023) with an 8B-parameter model, Llama 3 increases multilinguality, coding, reasoning, and tool usage Llama Team, Meta (2024). It offers a context length up to 128K, uses grouped-query attention for faster inference, and applies Direct Preference Optimization (DPO) and rejection sampling to align with human preferences, achieving competitive performance across multiple benchmarks.
@@ A.2 Proprietary Model GPT-4V
Gemini-1.5-Pro
With context length up to one million tokens, Gemini-1.5-Pro Gemini Team, Google (2024a) uses a Mixture-of-Experts (MoE) Shazeer et al. (2017) design for efficiency. Compared to Gemini 1.0 Gemini Team, Google (2024b), it achieves better multimodal reasoning (text, images, video, code), offers in-context learning, and integrates safety/ethics testing throughout development.
Appendix B Additional Experiments on Qwen-VL
To better understand multilingual and multicultural understanding abilities of our proposed pipeline, we performed additional experiments on Qwen-VL. This is because Qwen-VL is trained on Chinese data, while all other open models we evaluated are pretrained with a focus on English culture, without seeing much data from the local culture. Therefore, in addition to the experiments we discussed in Section 6.3, we also performed the third intervention with Qwen-VL on the MaRVL Mandarin Chinese dataset where we caption images using the native language. This experiment resulted in 73.4% accuracy, while using our interventions with English captions gives 72.8% accuracy, and using Qwen without interventions gives 58.4% accuracy. These results extended our points that visio-linguistic models need better understanding of culturally-specific elements. For example, Siheyuan is a culturally specific concept from Chinese culture, where if a model has never seen such concepts previously, it might not be able to generate the correct response for queries containing the concept Siheyuan.
Appendix C Additional Computation Cost
For the first intervention in §6.1, we use the translated statements provided in the MaRVL dataset, so no additional training cost is incurred.
For the second intervention in §6.2, training cost is not directly comparable, since we finetune ViLT if not using the intervention, and use the pretrained ViLT if using the intervention.
For the third intervention, with a 3% increase in total evaluation time, we see a 13% average improvement in performance for LLaVA-v1.5-13B. There is no additional training cost brought by the intervention. Noteworthily, total inference time using LLaVA is halved when using this intervention.
Appendix D Machine Translation V.S. Human Translation
In the translation test described in Section 6.1, we used the Google Translate API Wu et al. (2016). To investigate whether potential translation inaccuracies could omit certain linguistic nuances in non-English contexts, we also evaluated models on a human-translated version of the dataset.
Model Machine (zh) Human (zh)
xUNITER 63.3 64.4
GPT-4V 78.4 79.9
Table 5: MaRVL-ZH results (Machine translation vs. Human translation).
Specifically, we tested xUNITER (finetuned on NLVR2) and GPT-4V (zero-shot) on the Chinese subset of MaRVL using human-translated data provided by the original MaRVL paper Liu et al. (2021). Table 5 shows the results. Notably, the human-translated data yields only marginal improvements over machine translation. While we acknowledge the limitations inherent in the translation-based approach, these findings support using machine translation for broader evaluations due to its practicality under resource constraints.
Moreover, the small discrepancy in performance between human and machine translations suggests that the translation method itself may have only minor influences on model performance. Accordingly, we relied on the Google Translate API, consistent with the MaRVL translate test setting Liu et al. (2021).